A 20-lb bag cement mix contains 25% cement. How many pounds of pure cement must be added to the mixture in order to get 40% cement?
Answer by richwmiller(17219)

1*x+0.25*20=0.4(20+x)
1*x+5=8+0.4x
1x-0.4x=8-5
0.6x=3
x=5 lbs of 100% is added to 20 lbs of 25% to get 25 lbs at 40%
check
1*5+0.25*20=0.4(20+5)
5+5=0.4(25)
10=10
ok